### v1

# Saliency-Guided DETR for Moment Retrieval and Highlight Detection

Abstract
Existing approaches for video moment retrieval and highlight detection are not able to align text and video features efficiently, resulting in unsatisfying performance and limited production usage. To address this, we propose a novel architecture that utilizes recent foundational video models designed for such alignment. Combined with the introduced Saliency-Guided Cross Attention mechanism and a hybrid DETR architecture, our approach significantly enhances performance in both moment retrieval and highlight detection tasks. For even better improvement, we developed InterVid-MR, a large-scale and high-quality dataset for pretraining. Using it, our architecture achieves state-of-the-art results on the QVHighlights, Charades-STA and TACoS benchmarks. The proposed approach provides an efficient and scalable solution for both zero-shot and fine-tuning scenarios in video-language tasks. The dataset and code will be published as open-source [^1] https://github.com/ai-forever/sg-detr.
1 Introduction
The task of searching for specific moments in a video based on a text query has always been in demand. With the rapid growth of video content on all platforms, it has become increasingly important. At the same time, despite their improved quality, existing solutions still have very limited performance and, therefore, limited real-world application.
Existing methods mostly utilize frozen encoders, such as SlowFast [7] and CLIP [28], to extract features from videos and text queries in the QVHighlights[16] benchmark, but they often fail to align these features effectively. Some works, such as [22], have attempted to address CLIP’s limitations and enhance the quality of video features and have even achieved some performance improvement. However, recent advances in foundational video models have enabled the use of more powerful encoders for both text and video, paving the way for novel architectural approaches in the moment retrieval field. We propose a new model block that utilizes the advantages of the InternVideo2 [42] aligned video-text encoder. Our approach starts with generating initial highlight scores termed as Local Saliency Scores by directly comparing text and video embeddings and then refining them by incorporating the global context of the entire video.
Inspired by the success of DETR in 2D object detection [5], the authors of the QVHighlights benchmark adapted its principles for 1D video moment localization task. The resulting model, Moment-DETR [16], became a foundational development in the field, paving the way for further works [25, 24, 15, 36]. However, other architecture variants have also been proposed to address this problem. For instance, the authors of [19] avoid the one-to-one matching typical for DETR-based models, and instead tackle moment localization using standard detection heads with one-to-many matching. In our work, we combine both strategies. Inspired by advancements in 2D detection [49], we propose a hybrid architecture for detecting relevant intervals, which incorporates both a convolutional detector head and a transformer decoder block with a one-to-one matching scheme. Additionally, we use an IoU scoring mechanism to improve the localization of video spans further.
As was claimed in [24], not all video content may be relevant to the textual query. To mitigate this issue and incorporate relevant part of textual information into the video, we propose a Saliency-Guided Cross Attention mechanism for cross-modal interaction, where attention scores are weighted with Local Saliency Scores.
Another area of research in the field is the study of the interrelationship between the tasks of moment retrieval and highlight detection. Recently, the authors of TR-DETR [36] highlighted the need to ensure interaction between the heads solving these two tasks. We build upon these ideas and further integrate the tasks of moment retrieval and highlight detection using our Saliency Amplifier module.
Despite the high quality of the QVHighlights dataset’s annotations, its limited size necessitates additional data for effective model pretraining. To address this, we introduce a novel data generation approach that utilizes large video retrieval datasets, such as InterVid-FLT-10m [41], and leverages the impressive capabilities of foundational text-video models to create high-quality training data. This method enables us to build a pretraining dataset called InterVid-MR, which contains 150k samples. Our model, trained solely on the InterVid-MR dataset, demonstrates strong performance in zero-shot evaluations on the QVHighlights benchmark. Moreover, fine-tuned on QVHighlights data, the model achieves state-of-the-art results.
The contributions of our work can be summarized as:
  • We developed a new architecture for moment retrieval and highlight detection tasks, leveraging the latest advancements in video-text alignment and object detection.
  • We proposed a novel data generation approach and built a pre-training dataset with 150k samples.
  • We conducted extensive experiments to evaluate the effectiveness of the proposed model across three benchmarks: QVHighlights, Charades-STA and TACoS.
  • We provide publicly available training and inference code, InterVid-MR dataset and data generation scripts.
2 Related Works
Video Temporal Grounding
Video temporal grounding refers to linking a textual query to its corresponding segments within a video. This process involves two core components: moment retrieval and highlight detection.
Moment retrieval (MR) aims to localize specific video intervals based on a given textual query. Approaches for this task are typically classified into proposal-based and proposal-free methods. Proposal-based methods utilize predefined proposals, such as sliding windows [1, 10] or temporal anchors [6]. These methods often require complex pre- and post-processing, with performance being heavily influenced by the quality of the generated proposals. In contrast, proposal-free methods [16, 19, 24] leverage multimodal information and directly predict the start and end coordinates of the target moments, thus bypassing the need for explicit proposal generation.
Highlight detection (HD) focuses on assigning a relevance score to each video sub-clip based on the textual query, allowing for easy ranking. Methods in this domain are typically divided into three categories based on the level of annotation: supervised [12, 37], which requires precise and often expensive annotations; weakly supervised [4, 45], which learns to identify key moments using event labels; and unsupervised [2, 23, 32], which operates without the need for labeled data.
Traditionally, moment retrieval and highlight detection tasks have been addressed independently. However, a recent study [16] introduced the QVHighlights dataset, which facilitates the simultaneous resolution of both tasks within a unified framework. In conjunction with the dataset, the authors proposed a novel approach based on a DETR-like architecture designed to jointly tackle the challenges of both moment retrieval and highlight detection.
Building on the ideas proposed in [16], QD-DETR [25] enhanced video features by integrating textual information through Cross-Attention mechanisms. The subsequent study by [24] introduced an Adaptive Cross-Attention module that incorporates dummy tokens to redirect attention, thereby minimizing the representation of irrelevant video clips in response to the text query. BAM-DETR [15] extended DETR architecture by introducing a boundary-based prediction approach, where the model predicts segment boundaries rather than segment centers and their width. In [36], the authors explored the relationship between moment retrieval and highlight detection tasks and proposed a joint prediction framework that aligns both tasks.
Another line of research diverges from DETR-like architectures. UMT [21] introduced a transformer-based model that leverages multimodal data, integrating both video and audio streams to enhance performance. The most recent work, Mr.BLIP [3], proposed an approach based on LLMs, focusing exclusively on solving moment retrieval tasks.
Object Detection
The pioneering work that introduced a transformer-based object detector was DETR [5], which demonstrated the ability to predict objects end-to-end without relying on NMS or other post-processing techniques. Subsequent research has built upon this foundation, aiming to enhance performance and address the limitations inherent in the original approach.
Deformable DETR [48] addressed the slow convergence and high computational cost of the original DETR by introducing deformable attention. DN-DETR [17] employed a denoising approach, introducing noisy targets as queries to improve prediction accuracy and speed up convergence. DAB-DETR [20] further enhanced DETR by integrating dynamic anchor boxes into object queries, improving localization accuracy and guiding attention more effectively. Building on these advancements, DINO DETR [46] refined key features of both DN-DETR and DAB-DETR and integrated RPN into DETR architecture. CO-DETR [49] incorporated classical object detectors, such as ATSS [47] and Faster R-CNN [30], into the DETR framework, leveraging the strengths of both paradigms.
Multi-Task Learning (MTL)
Research in multimodal alignment, which integrates visual and textual modalities, has shown significant potential for improving performance in downstream tasks such as moment retrieval and highlight detection. Various methods have been proposed to tackle this challenge.
CLIP [28], SimVLM [43], and OFA [40], which operate in the image domain, have demonstrated strong multimodal learning capabilities by leveraging contrastive and generative learning methods. However, their application to video-based tasks like moment retrieval remains suboptimal because they primarily focus on static images. It constrains their ability to capture the temporal dynamics inherent in video content.
Recent studies have introduced models that are specifically designed for video understanding tasks. For example, InternVideo [41] utilized video masking and video-language contrastive learning as pretraining objectives, substantially improving video understanding. Further extending this approach, InternVideo2 [42] employed a multi-stage training strategy that integrated self-supervised and weakly supervised techniques, including masked video token reconstruction and cross-modal contrastive learning. InternVideo2 demonstrated significant improvements across various downstream tasks, especially in the moment retrieval domain.
Figure 1: Overall architecture of our framework. Detailed explanations of notations are described in Sec. 3
3 Model
Moment retrieval and highlight detection tasks aim to identify key moments in a video based on a textual query. In this section, we describe in detail each component of our proposed architecture, which effectively solves both tasks. An overview of the proposed Saliency-Guided DETR (SG-DETR) is shown in Figure 1.
3.1 Feature Extraction
Previous studies [16, 36, 25, 24] have used a combination of the CLIP [28] and SlowFast [7] models to extract features from the QVHighlights[16] dataset. They segmented the videos into 2-second intervals and extracted features from each segment using the CLIP and SlowFast models. These features were then concatenated.
This approach has several significant drawbacks. CLIP was initially trained to align images and texts. Thus, it does not fully account for the temporal dimension. This is why the SlowFast video extractor has to be added to the pipeline to compensate for the lack of spatial-temporal relationships in the CLIP features. However, the SlowFast and CLIP models were trained separately, so their embeddings are in different feature spaces.
Considering the issues mentioned above, we decided to use the InternVideoV2 [42] model, which was trained for aligning video and text. It effectively handles the modeling of spatial-temporal relationships and does not require the inclusion of additional models like SlowFast in the pipeline, significantly simplifying it.
Our feature extraction pipeline can be described as follows. We divide video into 2 seconds clips and extract features for each clip using the video encoder of the InternVideoV2-1B model: $F_{v}=[f_{v}^{1},f_{v}^{2},\ldots,f_{v}^{L}]\in\mathbb{R}^{L\times d_{v}}$, where $L$ and $d_{v}$ are the number of clips and the visual feature dimension, respectively. Similarly, text features are extracted using the text encoder of the InternVideoV2-1B model: $F_{t}=[f_{t}^{1},f_{t}^{2},\ldots,f_{t}^{N}]\in\mathbb{R}^{N\times d_{t}}$, where $N$ and $d_{t}$ are the number of the textual tokens and the textual feature dimension, respectively.
3.2 Local Saliency Scores
Features obtained from the InternVideo2 encoders [42] are already in a unified space, facilitating efficient and high-quality preliminary HD assessments. The scores obtained at this stage are termed as Local Saliency Scores $S_{\text{local}}$ because they do not take into account the global context of the entire video. To compute them, we first derive an overall vector of the text query using Pooling Encoder:
$F_{\text{sent}}=\text{PoolingEncoder}(F_{t}),$ (1)
where PoolingEncoder is an encoder consisting of $N$ layers of AttentionPooling modules, and $F_{\text{sent}}\in\mathbb{R}^{1\times d_{t}}$ represents the sentence token.
Next, we compute $S_{\text{local}}$ based on the similarity between the sentence token and the video embeddings. Formally, the local saliency scores are given by:
$S_{\text{local}}=\alpha\cdot\text{L2-Norm}(F_{\text{sent}})\cdot\text{L2-Norm}(F_{v})^{T}+\beta,$ (2)
where $\alpha$ and $\beta$ are learnable scaling and shifting factors, respectively. The use of L2 normalization and affine transformation helps to stabilize the training process. To make $F_{\text{sent}}$ token even more informative we apply moment-sentence alignment as described in [24].
3.3 Saliency-Guided Cross Attention
As it was highlighted in the CG-DETR paper [24], not all video content aligns with the semantics of a text query. However, when softmax is used as the key element of cross-attention, it can result in assigning similar importance to all query tokens even for irrelevant ones. This behavior can reduce the accuracy of the model in identifying the true correspondence between the video content and the text query. Although switching from softmax to sigmoid seems like a straightforward solution, it introduces limitations by removing the dependencies between tokens, which weakens the ability to rank text-relevance scores accurately.
To address these issues, we introduce the Saliency-Guided Cross Attention (SGCA) mechanism, which effectively overcomes the limitations of standard cross-attention by considering the local relevance of each individual video clip to the text query. Formally, let the query tokens be derived from the video tokens as $Q=[p_{Q}(f_{v}^{1}),p_{Q}(f_{v}^{2}),\ldots,p_{Q}(f_{v}^{L})]$, key and value tokens be derived from the text tokens: $K=[p_{K}(f_{t}^{1}),p_{K}(f_{t}^{2}),\ldots,p_{K}(f_{t}^{L})]$ and $V=[p_{V}(f_{t}^{1}),p_{V}(f_{t}^{2}),\ldots,p_{V}(f_{t}^{L})]$, where $p_{Q}(\cdot)$, $p_{K}(\cdot)$, and $p_{V}(\cdot)$ are learned projection functions that map the video and text tokens to a common hidden dimension $h$.
SGCA for the $i$-th video token $f_{v}^{i}$ is computed as:
$\text{SGCA}(f_{v}^{i})=\sum_{j=1}^{L}W_{i,j}\odot V_{j},$ (3)
where $W_{i,j}$ represents the attention weight for aligning the i-th video token with the j-th text token and computed as
$W_{i,j}=\text{Softmax}\left(\frac{Q_{i}\odot K_{j}}{\sqrt{h}}\right)\cdot\sigma(S_{\text{local}}^{j})$ (4)
The described attention mechanism is then used as a part of the Cross-Modal Transformer Encoder to derive the video token representation enriched with textual information.
${F_{v}}^{T}=\text{CrossModalEncoder}(F_{v},F_{t})$ (5)
After that, we apply a standard Transformer Encoder consisting of N layers to incorporate global context into features.
$\widehat{{F_{v}}^{T}}=\text{TransformerEncoder}({F_{v}}^{T})$ (6)
3.4 From Local to Global Saliency scores
The local saliency scores can be further refined based on $\widehat{{F_{v}}^{T}}$. To do that, we first incorporate the information from the computed local scores into $\widehat{{F_{v}}^{T}}$ using the Saliency Amplifier (SA) module, which can be described as follows:
$\widehat{{{F_{v}}^{T}}_{l}}=\widehat{{F_{v}}^{T}}+\widehat{{F_{v}}^{T}}\cdot\sigma(S_{\text{local}}),$ (7)
These features are used to compute the offsets for the local saliency scores to get the global saliency scores:
$S_{\text{global}}=S_{\text{local}}+\text{Linear}(\widehat{{{F_{v}}^{T}}_{l}}).$ (8)
We reapply the SA module once the global saliency scores are obtained. This time, the goal is to leverage the information obtained during the Highlight Detection task to address the Moment Retrieval task, as we now know which features are more relevant and which are less.
$\widehat{{{F_{v}}^{T}}_{g}}=\widehat{{F_{v}}^{T}}+\widehat{{F_{v}}^{T}}\cdot\sigma(S_{\text{global}}),$ (9)
We sent globally amplified video features $\widehat{{{F_{v}}^{T}}_{g}}$ to regression heads.
3.5 Hybrid Detector
In previous works, various detection heads, including CNN-based [19] and DETR-like models [25, 24, 16], have tackled the moment retrieval task. Co-DETR [49] has demonstrated that combining these approaches can be adequate for object detection tasks in the image domain. Inspired by this, we employ a hybrid detector with two detection heads. The ATSS [47] detector serves as the auxiliary CNN-based head, while the DINO-DETR [46] detector is used as the primary detection head. In line with Co-DETR, we utilize positive anchors from the ATSS detector as queries for the DETR head during model training. Consequently, during training, the DETR head receives three sets of references: a few groups of noisy target spans, positive anchors from ATSS, and primary references generated by the query selector mechanism introduced in DINO-DETR. As part of the post-processing stage, we follow [34] and apply a weighted fusion of ATSS and DETR boxes to produce the final predictions.
3.6 Localization confidence
Prior research, including works such as [14, 44, 9], has demonstrated that predicting IoU scores or a similar localization confidence metric significantly improves object detection performance. Building on these insights, we introduce an additional DETR head dedicated to predicting IoU scores between the predicted and ground truth spans.
We incorporate predicted IoU scores into the matching process. Additionally, predicting IoU scores enables a better ranking of the predicted spans. Similar to the work [44], we computed confidence of the span as a combination of classification and localization confidences:
$S_{det}=p_{i}^{\alpha}\hat{IoU_{i}}^{(1-\alpha)}$ (10)
where $p_{i}$ is the classification confidence and $\hat{IoU_{i}}$ is the predicted IoU.
3.7 Training Objectives
Our model is trained using losses categorized into three main groups.
Highlight Detection Task
We employ margin ranking, rank contrastive, and cross-entropy (CE) losses on both local and global saliency scores, as defined in [24]. Additionally, we apply the CE loss to negative text-video pairs following [25] to suppress negative clip saliency. The total loss for the task is represented as:
$\mathcal{L}_{hl}=\mathcal{L}_{marg}+\mathcal{L}_{rctl}+\mathcal{L}_{bce\_pos}+\mathcal{L}_{bce\_neg}$ (11)
Moment Retrieval Task
For the moment retrieval task, we use CE loss, generalized IoU (GIoU) loss [31], and smooth L1 loss to train both DETR and ATSS detection heads. For the ATSS head [47], we also incorporate Centerness Loss [39] to enhance localization precision. In the DETR head, CE loss is applied to IoU head predictions. Losses for auxiliary DETR queries are computed similarly to primary queries. The objectives for the task are:
$\begin{split}\mathcal{L}_{\text{detr}}&=\lambda_{\text{L1}}\mathcal{L}_{\text{L1}}(m,\bar{m})+\lambda_{\text{gIoU}}\mathcal{L}_{\text{gIoU}}(m,\bar{m})\\ &\quad+\lambda_{\text{CE}}\mathcal{L}_{\text{CE}}(y,\bar{y})+\lambda_{\text{IoU}}\mathcal{L}_{\text{CE}}(IoU,\bar{IoU})\end{split}$ (12)
$\begin{split}\mathcal{L}_{\text{atss}}&=\lambda_{\text{L1}}\mathcal{L}_{\text{L1}}(m,\bar{m})+\lambda_{\text{gIoU}}\mathcal{L}_{\text{gIoU}}(m,\bar{m})\\ &\quad+\lambda_{\text{CE}}\mathcal{L}_{\text{CE}}(y,\bar{y})+\lambda_{\text{Centrness}}\mathcal{L}_{\text{CE}}(c,\bar{c})\end{split}$ (13)
In this context, the ground truth values are represented as $m=(m_{c},m_{\sigma})$, $y$, $c$, and $IoU$, where $m_{c}$ and $m_{\sigma}$ denote the center and duration of the ground-truth moment, $y$ represents the binary classification label, $c$ is the target centerness score, and $IoU$ refers to the target IoU score. Similarly, the predicted values are denoted as $\bar{m}=(\bar{m}_{c},\bar{m}_{\sigma})$, $\bar{y}$, $\bar{c}$, and $\hat{IoU_{i}}$, corresponding to the predicted moment, binary classification label, centerness score, and predicted IoU score, respectively.
Auxiliary Losses
The third category includes auxiliary losses to enhance the model’s overall performance. First, alignment loss ensures consistency between moment and sentence token. Additionally, CE loss is applied to differentiate moment tokens from non-moment tokens within each video instance. Detailed descriptions of these losses can be found in [24]. The overall objective of the group is:
$\mathcal{L}_{aux}=\mathcal{L}_{bce}+\mathcal{L}_{align}$ (14)
Overall Objective
The final objective function is the sum of all the aforementioned losses:
$\mathcal{L}_{obj}=\lambda_{\text{atss}}\mathcal{L}_{\text{atss}}+\lambda_{\text{detr}}\mathcal{L}_{\text{detr}}+\lambda_{\text{hl}}\mathcal{L}_{hl}+\lambda_{\text{aux}}\mathcal{L}_{aux}$ (15)
4 Experiments and Results
Table 1: Comparison of models performance on QVHighlights val split using different feature extractors. Results are reported as mean $\pm$ standard deviation, averaged over three runs.
Method Features MR mAP HD $\geq$ Very Good
@0.5 @0.75 Avg. mAP HIT@1
Moment-DETR [16] CLIP+SlowFast - - 32.20 35.65 55.55
Moment-DETR InternVideo2-1b 60.20 $\pm$ 0.55 34.43 $\pm$ 0.43 35.40 $\pm$ 0.41 40.31 $\pm$ 0.21 63.89 $\pm$ 0.62
UniVTG [19] CLIP+SlowFast - - 36.13 38.83 61.81
UniVTG InternVideo2-1b 63.51 $\pm$ 0.25 38.83 $\pm$ 0.26 37.78 $\pm$ 0.16 42.68 $\pm$ 0.09 69.34 $\pm$ 0.23
QD-DETR [25] CLIP+SlowFast 62.23 41.82 41.22 39.13 63.03
QD-DETR InternVideo2-1b 67.78 $\pm$ 0.29 46.40 $\pm$ 0.26 45.52 $\pm$ 0.15 41.82 $\pm$ 0.07 68.06 $\pm$ 0.24
CG-DETR [24] CLIP+SlowFast 65.60 45.70 44.90 40.80 66.70
CG-DETR InternVideo2-1b 69.86 $\pm$ 0.21 49.35 $\pm$ 0.28 48.69 $\pm$ 0.17 42.72 $\pm$ 0.07 69.87 $\pm$ 0.15
TR-DETR [36] CLIP+SlowFast 66.27 46.42 45.09 40.55 64.77
TR-DETR InternVideo2-1b 70.08 $\pm$ 0.15 49.20 $\pm$ 0.50 47.99 $\pm$ 0.42 43.43 $\pm$ 0.16 71.13 $\pm$ 0.25
Ours InternVideo2-1b 73.52 $\pm$ 0.05 57.91 $\pm$ 0.13 55.64 $\pm$ 0.20 43.91 $\pm$ 0.14 71.47 $\pm$ 0.73
4.1 Datasets
We evaluate the effectiveness of the proposed method using three main benchmarks: QVHighlights [16], Charades-STA [8], TACoS [29].
QVHighlights is the primary benchmark annotated for both moment retrieval and highlight detection tasks. The dataset is relatively small and consists of approximately 10k videos with human-written text queries covering various topics, from everyday activities and travel in lifestyle vlog videos to social and political activities in news videos. Charades-STA is a dataset used exclusively for moment retrieval task evaluation. It includes 16k sentence-moment pairs from 10k videos of indoor activities with an average duration of 30 seconds. TACoS consists of 127 cooking videos with an average duration of 287 seconds and 19k sentence-moment pairs. This dataset is recognized as particularly challenging because the moments occupy only a tiny portion of the considerably long videos.
4.2 Feature Extraction
In earlier studies [21, 16, 25, 13], various feature extractors were used on the mentioned benchmarks. For the QVHighlights benchmark, a combination of CLIP [28] and SlowFast [7] models was employed. The Charades-STA benchmark used the VGG [33] extractor for video content and GLOVE [27] embeddings for text queries. This variety in methods requires maintaining numerous feature extraction pipelines and tuning hyperparameters for each, which slows down research and increases the carbon footprint. Additionally, many of these methods have become outdated and hardly remain representative. We propose using a unified approach to overcome the problem. Specifically, we extracted visual and textual features for each benchmark using only the InterVidV2-1b [42] model. We retrained the primary existing models using the InterVidV2-1b embeddings to assess the new features’ impact and validate our method’s competitiveness on QVHighlights validation set. The results are presented in Tab. 1.
4.3 Pretraining Framework
4.3.1 Motivation
The QVHighlights dataset contains only 10k training samples. Previous studies [16, 19] have shown that this dataset size is insufficient for training a model to effectively address a complex task like Moment Retrieval. As a result, even pretraining on a noisy dataset can significantly improve the model’s metrics. For instance, pretraining framework developed in [19] boosted the model performance from 35.47 to 43.63 mAP@avg. Nevertheless, the metrics in the zero-shot mode remained relatively low at only 10.87 mAP@avg [19], leading us to believe that creating a dataset for pretraining remains an unsolved challenge.
4.3.2 InterVid-MR Dataset
Recent advancements in video-language modeling have been driven by the annotation of large scale, high-quality datasets like InternVid [41]. Unlike VideoCC [26] and other datasets previously used for pretraining [11, 16], InternVid was annotated using LLMs and does not rely on pseudo-labeling, resulting in significantly less annotation noise. To construct our dataset, we used the subset InternVid-10M-FLT [41] derived from the original InternVid by sampling clips with UMT-SIM scores [18] ranking among the top 30 percent to ensure high quality.
Unfortunately, InternVid-10M-FLT cannot be directly used to train MR models because each caption is linked to only one video segment in the annotations, whereas in practice multiple video segments may closely match a given query. We have made several steps to address this issue and developed a new dataset named InterVid-MR, which is suitable for Moment Retrieval and Highlight Detection tasks.
First, we additionally filtered the captions of the InternVid-10M-FLT dataset to remove meaningless ones. Next, to align the data with the QVHighlights dataset, we selected video segments no longer than 150 seconds for each caption from the InternVid-10M-FLT dataset. These segments were created by trimming the original videos around the time intervals associated with the captions.
Then, we split each video into 2-second clips and extracted features using the IntervidV2 video encoder. Caption features were extracted using the IntervidV2 text encoder. We computed the cosine similarity between the embeddings of each video clip and the corresponding caption embedding using the formula:
$S=\text{L2\_Norm}(F_{v})\cdot\text{L2\_Norm}(F_{c})^{T},$ (16)
where $F_{v}$ represents the embedding of the video clip, and $F_{c}$ represents the caption embedding.
We used the similarity scores of video clips within the original positive interval, defined by the dataset annotation, to determine the minimum similarity score $S_{\text{min}}$ required for a clip to be included in a new positive interval. This minimum similarity score $S_{\text{min}}$ is calculated as:
$S_{\text{min}}=\operatorname{mean}(S_{\text{pos}})-3\times\operatorname{std}(S_{\text{pos}})$ (17)
where $S_{\text{pos}}$ represents the similarity scores of all clips within the original positive interval. Based on the $S_{\text{min}}$ threshold, we constructed new positive intervals consisting of clips whose similarity scores are higher than $S_{\text{min}}$.
After annotating the Moment Retrieval task, we proceeded to annotate the Highlight Detection task. We used the similarity scores calculated in the previous step to assign each clip within the identified positive intervals a grade ranging from 1 to 4. The grades are calculated using the following formula:
$\text{Grade}(S)=\begin{cases}4,&\text{if }S\in(\mu+1.5\times\sigma,S_{\text{max}}]\\ 3,&\text{if }S\in(\mu,\mu+1.5\times\sigma]\\ 2,&\text{if }S\in(\mu-1.5\times\sigma,\mu]\\ 1,&\text{if }S\in[S_{\text{min}},\mu-1.5\times\sigma]\end{cases}$ (18)
where $S$ is the score from the positive interval to be graded, $\mu$ is the mean score of the positive clips, $\sigma$ is the standard deviation of the positive clips’ scores, and $S_{\text{max}}$ is the maximum score of the positive clips.
As a result, we obtain a dataset that can be used to solve both Moment Retrieval and Highlight Detection tasks simultaneously.
4.3.3 Impact of the Pretraining
Using our data annotation framework, we collected a dataset consisting of 150k samples. To ensure the effectiveness of the proposed data annotation strategy for solving Moment Retrieval and Highlight Detection tasks, we measured the zero-shot performance of a model trained on different amounts of pre-train data on the QVHighlights dataset. As shown in Figure  2, even 20k samples in the training dataset are sufficient to achieve an MR mAP@avg above 45. Using all the data available, we surpassed the 50 MR mAP@avg threshold, which significantly exceeds the results of the pretraining strategy proposed in [19].
30303030606060609090909012012012012015015015015045454545505050505555555560606060Number of Pretrain PointsmAP@avg (%)QVHighlightsPretrainFinetune
Figure 2: The impact of pre-train dataset size on MR mAP@avg metric on the QVHighlights validation set.
Table 2: Ablation study on QVHighlights val split. SGCA and GSM stand for saliency-guided cross-attention transformer encoder and global saliency module, respectively. Results are reported as mean $\pm$ standard deviation, averaged over three runs.
MR HD
SGCA IoU Scoring GSM Hybrid DETR R1 mAP $\geq$ Very Good
@0.5 @0.7 @0.5 @0.75 Avg. mAP HIT@1
(a) 70.1 $\pm$ 0.3 53.7 $\pm$ 0.5 70.0 $\pm$ 0.1 49.3 $\pm$ 0.2 48.3 $\pm$ 0.4 41.9 $\pm$ 0.1 68.6 $\pm$ 0.8
(b) 71.1 $\pm$ 0.5 56.5 $\pm$ 0.8 71.0 $\pm$ 0.4 52.2 $\pm$ 0.7 50.1 $\pm$ 0.6 42.2 $\pm$ 0.1 70.1 $\pm$ 0.6
(c) 70.8 $\pm$ 0.2 56.5 $\pm$ 0.5 71.3 $\pm$ 0.3 52.5 $\pm$ 0.3 51.1 $\pm$ 0.2 42.3 $\pm$ 0.1 69.9 $\pm$ 0.5
(d) 72.1 $\pm$ 0.9 57.6 $\pm$ 0.4 72.6 $\pm$ 0.3 53.6 $\pm$ 0.5 52.2 $\pm$ 0.3 43.8 $\pm$ 0.2 71.1 $\pm$ 0.4
(e) 72.8 $\pm$ 0.5 59.5 $\pm$ 0.2 73.5 $\pm$ 0.1 57.9 $\pm$ 0.1 55.6 $\pm$ 0.2 43.9 $\pm$ 0.1 71.4 $\pm$ 0.7
4.4 Training Details
We train SG-DETR using two attention-pooling layers in the Local Saliency Head and two cross-attention transformer layers in the SGCA Encoder. The model architecture includes three layers of encoders and three layers of decoders. The hidden dimension of the transformer layers is set to 256. we employ the AdamW optimizer with a weight decay of 1e-1. A dropout rate of 0.1 is applied throughout the model, with a higher dropout rate of 0.5 in the projection layers. Our query selector mechanism generates 25 queries per instance. Unless otherwise specified, the following loss balancing parameters are used: $\lambda_{\text{L1}}=10$, $\lambda_{\text{gIoU}}=1$, $\lambda_{\text{CE}}=5$, $\lambda_{\text{Centerness}}=1$, $\lambda_{\text{detr}}=1$, $\lambda_{\text{atss}}=1$, $\lambda_{\text{hl}}=1$, $\lambda_{\text{aux}}=1$.
Dataset-Specific Settings:
  • QVHighlight: The model is trained for 150 epochs with a batch size of 128 and a learning rate of 5e-4. We employ the WarmupMultiStepLR scheduler, utilizing 45 warmup steps, with two milestones set at 100 and 125 epochs. The learning rate is decayed by a factor of 0.5 at each milestone.
  • Charades-STA: The model is trained for 100 epochs with a batch size of 64 and a learning rate of 5e-4. Video features are extracted from 1-second clips. During training, binary highlight scores were generated based on whether a clip overlapped more than halfway with a relevant moment retrieval interval
  • TACoS: The model is trained for 120 epochs with a batch size of 32 and a learning rate of 5e-4. Video features are extracted from 2-second clips. Binary highlight scores were generated the same way as in Charades-STA.
Table 3: Experimental results on the QVHighlights test set. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training. † denotes using audio modality. Bold letters indicate the best results.
MR HD
Method R1 mAP $\geq$ Very Good
@0.5 @0.7 @0.5 @0.75 Avg. mAP HIT@1
MDETR [16] 52.9 33.0 54.8 29.4 30.7 35.7 55.6
UMT † [21] 56.2 41.2 53.4 37.0 36.1 38.2 60.0
UniVTG [19] 58.9 40.9 57.6 35.6 35.5 38.2 61.0
QD-DETR [25] 62.4 45.0 62.6 39.9 39.9 38.6 62.4
CG-DETR [24] 65.4 48.4 64.5 42.8 42.9 40.3 66.2
BAM-DETR [15] 62.7 48.6 64.6 46.3 45.4 - -
TR-DETR [36] 64.7 49.0 64.0 43.7 42.6 39.9 63.4
Mr. BLIP [3] 74.7 60.5 68.1 53.4 51.4 - -
Ours 72.2 56.6 73.2 55.8 54.1 43.76 69.13
Ours ZS 63.9 49.6 67.5 49.0 48.3 43.0 68.0
Ours w/ PT 74.2 60.4 76.2 60.8 58.8 44.7 71.0
Table 4: Experimental results on the Charades-STA and TACoS test sets. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training. † denotes using audio modality. Bold letters indicate the best results.
Charades-STA TACoS
Method R@0.5 R@0.7 mIoU R@0.5 R@0.7 mIoU
MDETR [16] 53.6 31.4 - 24.7 12.0 25.5
UMT † [21] 48.3 29.3 - - - -
UniVTG [19] 58.0 35.7 50.1 35.0 17.4 33.6
QD-DETR [25] 57.3 32.6 - 36.8 21.1 35.8
CG-DETR [24] 58.4 36.3 50.1 39.6 22.2 36.5
BAM-DETR [15] 60.0 39.4 52.3 41.5 26.8 39.3
TR-DETR [36] 57.6 33.5 - - - -
MR. BLIP [3] 69.3 49.3 58.6 - - -
Ours 70.2 49.5 59.1 44.7 29.9 40.9
Ours w/ PT 71.1 52.8 60.7 46.4 33.9 42.4
4.5 Comparison with the state-of-the-art
Joint Moment Retrieval and Highlight Detection
To evaluate the effectiveness of SG-DETR in jointly addressing Moment Retrieval and Highlight Detection tasks, we conducted experiments on the QVHighlights test split, as shown in Table 3. Remarkably, even without pretraining, SG-DETR significantly outperforms all existing methods that do not leverage LLMs, and it delivers performance comparable to the MrBLIP model despite the latter having considerably more parameters. After pretraining on our constructed InterVid-MR dataset, SG-DETR achieves competitive zero-shot performance, with 48.3 mAP and 68.0 HIT@1 results. When fine-tuned on the QVHighlights dataset, the pre-trained SG-DETR surpasses all existing methods, including those based on LLMs, with 58.8 mAP and 71.0 HIT@1 results.
Moment Retrieval
We use the Charades-STA and TACoS datasets as benchmarks to evaluate the performance of SG-DETR exclusively in the moment retrieval task. The results are presented in Table 4. SG-DETR achieves state-of-the-art performance on these benchmarks, both with and without leveraging additional data from our pretraining framework. The results highlight the model’s robustness and ability to generalize effectively across varying task conditions.
4.6 Ablation study
To evaluate the contribution of each component in our approach, we conducted a comprehensive ablation study, the results of which are shown in Table 2. The baseline model (a) consists of a DETR-based architecture [16, 19], augmented with the local saliency head described in Sec. 3.2 and employing standard cross-attention as the text-to-vision interaction mechanism [25]. In experiment (b), we assess the effect of incorporating the SGCA encoder, as detailed in Sec. 3.3. It increases the speed of convergence of the model and pushes model performance forward, especially the moment retrieval part. Experiment (c) introduces the IoU scoring mechanism, described in Sec. 3.6, resulting in further growth of moment retrieval performance. Next, in experiment (d), we investigate the influence of the GSM module, discussed in Sec. 3.4 on model performance. It sufficiently improves not only moment retrieval performance but especially the quality of HD. Finally, in experiment (e), we replace the standard DETR head with a hybrid ATSS-DETR head, as outlined in Sec. 3.5, drastically improving MR performance.
5 Conclusions
We introduced the SG-DETR, a unified framework for Moment Retrieval and Highlight Detection, and presented the new InterVid-MR dataset for pretraining. Evaluations on benchmarks such as QVHighlights, Charades-STA and TACoS demonstrate that the SG-DETR sets new state-of-the-art results across all tasks. Our ablation studies highlight the effectiveness of key components like the Saliency-Guided Cross-Attention encoder and hybrid ATSS-DETR head, offering a scalable and robust solution for video-language modeling.
6 Limitations
Moment Retrieval and Highlight Detection tasks present significant challenges due to the ambiguity of textual queries and the subjective nature of video content. While SG-DETR achieves state-of-the-art performance, it may face difficulties with abstract or nuanced queries where the boundaries of relevant video segments are less clearly defined.
7 Acknowledgements
We want to express our sincere gratitude to Mikhail Chernyshov for his valuable insights and constructive discussions, which significantly contributed to shaping the direction of the work.
Supplementary Material
A Results on Highlight Detection Benchmarks
To evaluate the effectiveness of SG-DETR in addressing the highlight detection task independently, we evaluate its performance on the Youtube-hl [38] and TVSum [35] benchmarks. YouTube Highlights consists of 433 videos divided into six categories, and TVSum consists of 50 videos divided into ten categories. The results are presented in Table 5 and Table 6. The SG-DETR model achieves state-of-the-art performance on the YouTube-HL benchmark when additional data is utilized. Furthermore, SG-DETR achieves highly competitive results even in the zero-shot setting. The TVSum dataset contains only a limited number of videos for training in each domain, which appears to be insufficient for our model. Although the model achieves state-of-the-art performance specifically in the VT, PK, and BT domains, it overall performs slightly worse than TR-DETR. Additionally, due to the instability of training on such a small dataset, fine-tuning the model with additional data does not yield any improvements.
Table 5: Experimental results of mAP on the Youtube-hl. † denotes using audio modality. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training, Bold letters indicate the best results.
Method Dog Gym. Par. Ska. Ski. Sur. Avg.
UMT† [21] 65.9 75.2 81.6 71.8 72.3 82.7 74.9
UniVTG [19] 71.8 76.5 73.9 73.3 73.2 82.2 75.2
QD-DETR [25] 72.2 77.4 71.0 72.7 72.8 80.6 74.4
CG-DETR [24] 76.3 76.1 76.0 75.1 81.9 75.9 75.9
Ours 75.1 79.7 76.9 71.0 74.5 83.2 76.7
Ours ZS 65.0 69.1 76.6 57.5 76.1 70.8 69.2
Ours w/ PT 66.7 78.3 86.5 78.6 75.9 82.2 78.0
Table 6: Experimental results of Top-5 mAP on the TVSum. † denotes using audio modality. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training, Bold letters indicate the best results.
Method VT VU GA MS PK PR FM BK BT DS Avg.
UMT† [21] 87.5 81.5 88.2 78.8 81.5 87.0 76.0 86.9 84.4 79.6 83.1
UniVTG [19] 83.9 85.1 89.0 80.1 84.6 81.4 70.9 91.7 73.5 69.3 81.0
QD-DETR [25] 88.2 87.4 85.6 85.0 85.8 86.9 76.4 91.3 89.2 73.7 85.0
CG-DETR [24] 86.9 88.8 94.8 87.7 86.7 89.6 74.8 93.3 89.2 75.9 86.8
TR-DETR [36] 89.3 93.0 94.3 85.1 88.0 88.6 80.4 91.3 89.5 81.6 88.1
Ours 89.9 90.5 88.0 87.4 90.0 84.5 77.9 93.2 89.9 80.3 87.1
Ours ZS 44.3 70.9 85.5 65.7 61.1 49.0 44.3 67.3 83.2 48.8 62.0
Ours w/ PT 86.6 74.2 85.9 82.9 90.6 74.5 69.0 92.6 87.6 67.7 81.2